# Extended Galerkin neural network approximation of singular variational problems with error control

### v1

@@ 3 Applications
We consider several indefinite and non-self adjoint boundary value problems and present well-posed least squares variational formulations for each of them. All hyperparameters for each example are provided in Appendix B. Of particular interest to us are problems whose solutions contain singular features, for instance due to low-regularity data or nonconvex domain corners.
@@
for all $(\mathbf{v},q)\in\mathbf{H^{1}_{0}}(\Omega)\times(L^{2}(\Omega)/\mathbb{R})$. Formulation (33) is not SPD and thus we must use (13) in conjunction with Galerkin neural networks. However, if $\Omega$ contains nonconvex corners, then $\mathbf{u}\notin\mathbf{H^{2}}(\Omega)$ and $p\notin H^{1}(\Omega)$, and the spaces $\mathcal{X}$, $\mathcal{V}$, and $\mathcal{W}$ must be chosen carefully to account for the reduced regularity.
@@ 3.1 Poisson equation
Figure 4: The true solution $u(r,\theta)=r^{\lambda}\sin{\theta}$ (left) and approximate solutions obtain using Galerkin neural networks and (36) with $\beta=0$ (middle) and $\beta=1$ (right).
Figure 5: Example 3.2: for $\beta=0$, the errors $u-u_{i-1}$ and learned basis functions $\varphi_{i}^{NN}$.
@@ 3.1.1 Regularity and variational formulation
The following a priori estimate is a result of Kondrat’ev [11].
@@
The spaces $H^{s}_{\beta}(\Omega)$ are weighted Sobolev spaces defined by
[…]
where $r_{i}$ is the local radial coordinate centered at vertex $x^{(i)}$. In what follows, we will commonly use $(\cdot,\cdot)_{\beta,\Omega}$ to denote the $L^{2}_{\beta}(\Omega)$ inner product. We propose the weighted least squares variational formulation given by
$u\in H^{2}_{\beta}(\Omega)\;:\;(\Delta u,\Delta v)_{\beta,\Omega}$ $+(u,v)_{H_{\beta}^{3/2}(\partial\Omega)}=-(f,\Delta v)_{\beta,\Omega}+(g,v)_{H_{\beta}^{3/2}(\partial\Omega)}$
@@
Figure 6: Example 3.2: for $\beta=1$, the errors $u-u_{i-1}$, learned basis functions $\varphi_{i}^{NN}$, and $L^{2}$- and energy norm errors.
@@
We consider the case of the true solution $u(r,\theta)=r^{\lambda}\sin{\theta}$ for $\lambda=1/4$ in the domain $\Omega=(-1,1)^{2}\backslash(-1,0)^{2}$. In this case, $f\notin L^{2}(\Omega)$ but $f\in L^{2}_{\beta}(\Omega)$ for $\beta>1-\lambda$. We demonstrate that if $\beta=0$ in (36), then the approximation obtained is incorrect and does not converge to the true solution. Figure 4 shows the approximation using 8 basis functions.
We observe spurious oscillations of the approximation near $x=0$ when $\beta=0$ which are largely removed when $\beta=1$. Figures 5-6 show the true errors $u-u_{i-1}$ and basis functions $\varphi_{i}^{NN}$ for $\beta=0$ and $\beta=1$, respectively. We observe that in the $\beta=0$ case, the learned basis functions contain many high frequency modes not present in the corresponding error functions. In the $\beta=1$ case, the basis functions take on large values near $x=0$ and correctly identify the presence of error the approximation of the singularity, but convergence is still slow.
Figure 7: Example 3.3: for $\beta=4/3$, the approximate solutions $u_{i}$ and learned basis functions $\varphi_{i}^{NN}$.
@@ 3.1.2 Knowledge-based functions
Suppose $\Omega$ contains a single corner of interest $x^{(0)}$ with interior angle $\alpha^{(0)}$. For instance, the $L$-shaped domain $(-1,1)^{2}\backslash(-1,0)^{2}$ contains a nonconvex corner with angle $\alpha^{(0)}=3\pi/2$ at the origin. Solutions of (34) in such a domain have the well-known structure
@@
$\lambda_{j}=\frac{j\pi}{\alpha^{(0)}},\;\;\;j=1,2,3,\dots.$
Figure 8: Example 3.3: for $\beta=4/3$ and singular neural network architecture, the approximate solutions $u_{i}$ and learned basis functions $\varphi_{i}^{NN}$.
@@
While there is no closed form solution to this problem, we provide the value of the a posteriori error indicator $\langle r(u_{i-1}),\varphi_{i}^{NN}\rangle$ for each basis function learned as well as the basis functions $\varphi_{i}^{NN}$ themselves, which provide a function representation of the error $u-u_{i-1}$. Figure 7 shows the first several basis functions using a standard feedforward neural network for learning each basis function. We observe that the approximation converges quite slowly near $x=0$ and the basis functions $\varphi_{i}^{NN}$ show that the magnitude of the pointwise error is large in the vicinity of the singularity.
Next, we repeat the simulation, this time with an augmented neural network architecture. Namely, we compute $\varphi_{i}^{NN}$ according to (23) with $\Phi(r,\theta;\lambda)=r^{\lambda}\sin(\lambda\theta)$. Since the eigenvalues are known, we simply take $\lambda_{j}=2j/3$ and $n_{\lambda}=20$. However, we shall demonstrate in Section 4 how Galerkin neural networks may be used to approximate the values of $\lambda$ when they are unknown. Figure 8 shows the analogous results to Figure 7. We observe significantly faster convergence and the approximate pointwise errors decrease steadily from $\mathcal{O}(10^{-2})$ to $\mathcal{O}(10^{-5})$. Additionally, we note that the singularity is clearly captured correctly by the knowledge-based functions.
@@ 3.2.1 Regularity and variational formulation
Most notable is that (3.4) suggests that the residual of the incompressibility equation in (32) should actually be taken in the $H^{1}_{\beta}$ norm, which leads to the continuous and coercive formulation: seek $(\mathbf{u},p)\in\mathbf{H^{2}_{\beta}}(\Omega)\times H^{1}_{\beta}(\Omega)/\mathbb{R}$ such that
$(-\Delta\mathbf{u}+\nabla p,-$ $\Delta\mathbf{v}+\nabla q)_{\beta,\Omega}+(\text{div}\;\mathbf{u},\text{div}\;\mathbf{v})_{H^{1}_{\beta}(\Omega)}+(\mathbf{u},\mathbf{v})_{H^{3/2}_{\beta}(\partial\Omega)}$
$=(\mathbf{f},-\Delta\mathbf{v}+\nabla q)_{\beta,\Omega}+(g,\text{div}\;\mathbf{v})_{H^{1}_{\beta}(\Omega)}+(\mathbf{u}_{D},\mathbf{v})_{\mathbf{H^{3/2}_{\beta}}(\Omega)}$
for all $(\mathbf{v},q)\in\mathbf{H^{2}_{\beta}}(\Omega)\times H^{1}_{\beta}(\Omega)/\mathbb{R}$. For ease of notation, we set $X:=\mathbf{H^{2}_{\beta}}(\Omega)\times H^{1}_{\beta}(\Omega)/\mathbb{R}$ endowed with the norm $||(\mathbf{v},q)||_{X}:=(||\mathbf{v}||_{H^{2}_{\beta}(\Omega)}^{2}+||q||_{H^{1}_{\beta}(\Omega)}^{2})^{1/2}$ and define the bilinear operator $B:X\times X\to\mathbb{R}$ by
(39) $B((\mathbf{u},p),(\mathbf{v},q)):=(-\Delta\mathbf{u}$ $+\nabla p,-\Delta\mathbf{v}+\nabla q)_{\beta,\Omega}$
$+(\text{div}\;\mathbf{u},\text{div}\;\mathbf{v})_{H^{1}_{\beta}(\Omega)}+(\mathbf{u},\mathbf{v})_{H^{3/2}_{\beta}(\partial\Omega)}$
and the linear operator $L:X\to\mathbb{R}$ by
(40) $L((\mathbf{v},q)):=(\mathbf{f},-\Delta\mathbf{v}+\nabla q)_{\beta,\Omega}+(g,\text{div}\;\mathbf{v})_{H^{1}_{\beta}(\Omega)}+(\mathbf{u}_{D},\mathbf{v})_{\mathbf{H^{3/2}_{\beta}}(\Omega)}.$
@@ 3.2.2 Knowledge-based functions
with the constants $A,B,C,D$ chosen so that $\Phi$ satisfies suitable boundary conditions.
Figure 9: Channel with triangular cavity for Example 3.5. The rectangle channel is described by $(-2,2)\times(0,2)$ while the triangular cavity has vertices at $(-1,0)$, $(1,0)$, and $(-3,0)$.
@@
We specify $\mathbf{f}=\mathbf{0}$ and $g=0$, in which case $2\text{div}\;\epsilon(\mathbf{u})=\Delta\mathbf{u}$. At the reentrant corners, we expect a singularity to manifest in both the velocity and pressure. At the bottom corner, we expect infinitely cascading eddies driven by the channel flow. As described in Section 3.2.2, we shall supplement the standard feedforward network with $\Psi(r,\theta;\lambda)=\chi(r)\mathfrak{Re}[r^{\lambda}\Psi_{\lambda}(\theta)]$ in the bottom corner and $\Psi(r,\theta;\lambda)=\chi(r)r^{\lambda}\Psi_{\lambda}(\theta)$ at the reentrant corners. The exact values of $\lambda$ are $\lambda=1.58223$ in the reentrant corners and $\lambda=7.56813+3.37941i$ in the bottom corner. Precise details may be found in Appendix A. As with Example 3.3, we shall demonstrate in Section 4 how these values of $\lambda$ may be approximated using Galerkin neural networks in the event that they are unknown.
Figure 10: Example 3.5: velocity and pressure with $\beta=0$ (top) and $\beta=5/3$ (bottom) in (39).
@@
We first demonstrate the importance of choosing the Sobolev weight parameter $\beta$ appropriately. In contrast with Example 3.2, this example contains smooth data $f$ and $g$ and continuous data $\mathbf{u}_{D}$. Figure 10 shows the approximated velocities and pressures for $\beta=0$. We shall use $\varphi_{1,i}^{NN}$, $\varphi_{2,i}^{NN}$, and $q_{i}^{NN}$ to denote the $i$th basis functions for the velocities and pressure, respectively. The resulting velocity and pressure fields in Figure 10 show that even after learning six basis functions, the flow is nonphysical; we expect a near constant velocity $u_{1}$ along the $x$-direction in the channel. Figure 10 also shows the analogous results for $\beta=5/3$. Despite not adequately capturing the singularities and the expected sequence of eddies, the approximate solutions do not exhibit spurious oscillations and the behavior in the channel is as expected.
@@
Figure 11: Example 3.5: velocity and pressure with $\beta=5/3$ in (39) and basis functions learned from extended neural network with knowledge-based functions.
@@
We fix $\beta=5/3$ next and approximate the solution using six basis functions learned by extended Galerkin neural networks. We remarkably observe that with the supplemental activation functions, i.e. $\varphi_{1,i}^{NN},\varphi_{2,i}^{NN},q_{i}^{NN}\in V_{\mathbf{n},L}^{\sigma}\oplus V_{n_{\lambda}}^{\Phi}$, both the singular features at the reentrant corners and the eddies in the cavity are captured accurately. In fact, as Figure 11 shows, we are able to capture infinitely many eddies. The basis functions used to approximate the velocities and pressures are shown in Figure 12.
@@
Finally, when $\beta$ is too large, the solution space contains functions which may have worse singular behavior than the true solution due to the strong damping effect of $r^{\beta}$ on the singularities. In particular, the bilinear operator $B$ may admit as arguments functions with far stronger singularities than those of the true solution. We show the results for $\beta=4$ in Figure 13.

### v2

@@ 3 Applications
We consider several indefinite and non-self adjoint boundary value problems and present (1) well-posed least squares variational formulations and (2) numerical results demonstrating the efficacy of the extended Galerkin neural network approach for each example. All hyperparameters for each example are provided in Appendix B. Of particular interest to us are problems whose solutions contain singular features, for instance due to low-regularity data or non-convex domain corners.
@@
for all $(\mathbf{v},q)\in\mathbf{H^{1}_{0}}(\Omega)\times(L^{2}(\Omega)/\mathbb{R})$. Formulation (48) is not SPD and therefore the basic Galerkin neural network formulation outlined in Section 2.1 is not applicable. Instead, one could attempt to use a least squares formulation based directly on (47). However, if $\Omega$ contains non-convex corners, then $\mathbf{u}\notin\mathbf{H^{2}}(\Omega)$ and $p\notin H^{1}(\Omega)$, and a pure $\mathbf{L}^{2}(\Omega)$-based least squares formulation would not be valid. The spaces $\mathcal{X}$, $\mathcal{V}$, and $\mathcal{W}$ must be chosen carefully to account for the reduced regularity. The same objection is valid for the Poisson equation. Fortunately, this issue can be resolved by choosing the spaces $\mathcal{X}$, $\mathcal{V}$, and $\mathcal{W}$ to be appropriately weighted Sobolev spaces. This approach is explained in the remainder of this section. For simplicity, we begin by considering the Poisson equation.
@@ 3.1.1 Regularity and variational formulation
As mentioned in Section 2, the singular solution structure for (49) takes the form $\Psi(r,\theta)=\sum_{\lambda}c_{\lambda}r^{\lambda}\sin(\lambda\theta)$, where $\lambda>1/2$. Similar to the Stokes problem, a pure $L^{2}$-based least squares formulation is not valid. However, the solution does belong to a weighted space $H^{2}_{\beta}(\Omega)$ provided that the weight $\beta>0$ is sufficiently large.
For $\beta=(\beta_{1},\dots,\beta_{M})\geqslant\mathbf{0}$ and $s>0$, the spaces $H^{s}_{\beta}(\Omega)$ are weighted Sobolev spaces equipped with the norm
(50) $||u||_{H^{s}_{\beta}(\Omega)}^{2}=||u||_{H^{s-1}(\Omega)}^{2}+\sum_{|\alpha|=s}||\prod_{i=1}^{M}r_{i}^{\beta_{i}}D^{\alpha}u||_{L^{2}(\Omega)}^{2},$
where $r_{i}$ is the local radial coordinate centered at vertex $x^{(i)}$. If $s=0$, then the term involving the $H^{s-1}(\Omega)$ norm is dropped. In what follows, we will commonly use $(\cdot,\cdot)_{\beta,\Omega}$ to denote the $L^{2}_{\beta}(\Omega)$ inner product. The corresponding spaces $\mathbf{H}^{s}_{\beta}$ and $\mathbf{L}^{2}_{\beta}$ of vector-valued functions are defined in the usual way.
A simple calculation shows that $r^{\lambda}\sin(\lambda\theta)\in H^{2}_{\beta}(\Omega)$ if and only if $\beta>1-\lambda$. Thus, one might expect a continuous dependence result of the form (12) to hold in the space $\mathcal{X}=H^{2}_{\beta}(\Omega)$, $\beta>1-\lambda$. The following result, proved in [12], confirms this expectation.
Figure 4: The true solution $u(r,\theta)=r^{\lambda}\sin{\theta}$ (left) and approximate solutions obtain using Galerkin neural networks applied to the variational problem (51) with $\beta=0$ (middle) and $\beta=1$ (right).
Figure 5: Example 3.2: for $\beta=0$, the exact errors $u-u_{i-1}$ (top row) and approximate errors which correspond to basis functions $\varphi_{i}^{NN}\approx u-u_{i-1}$ learned using the Galerkin neural network approach (bottom row).
@@
We propose the weighted least squares variational formulation given by
$u\in H^{2}_{\beta}(\Omega)\;:\;(\Delta u,\Delta v)_{\beta,\Omega}$ $+\delta(u,v)_{H_{\beta}^{3/2}(\partial\Omega)}=-(f,\Delta v)_{\beta,\Omega}+\delta(g,v)_{H_{\beta}^{3/2}(\partial\Omega)}$
@@
Figure 6: Example 3.2: for $\beta=1$, the exact errors $u-u_{i-1}$ (top row) and approximate errors which correspond to learned basis functions $\varphi_{i}^{NN}\approx u-u_{i-1}$ (middle row). The energy and and $L^{2}$ errors are provided for both $\beta=0$ and $\beta=1$.
@@
We consider the case of the true solution $u(r,\theta)=r^{\lambda}\sin{\theta}$ for $\lambda=1/4$ in the domain $\Omega=(-1,1)^{2}\backslash(-1,0)^{2}$. In this case, $f\notin L^{2}(\Omega)$ but $f\in L^{2}_{\beta}(\Omega)$ for $\beta>1-\lambda$. We demonstrate that if $\beta=0$ in (51), then the approximation obtained is incorrect and does not converge to the true solution. Figure 4 shows the approximation using 7 basis functions. All hyperparameters for this example are provided in Appendix B.
We observe spurious oscillations of the approximation near $x=0$ when $\beta=0$ which are largely removed when $\beta=1$. Figures 5-6 show the true errors $u-u_{i-1}$ and basis functions $\varphi_{i}^{NN}$ for $\beta=0$ and $\beta=1$, respectively. We observe that in the $\beta=0$ case, the learned basis functions contain many high frequency modes not present in the corresponding error functions. In the $\beta=1$ case, the basis functions take on large values near $x=0$ and correctly identify the presence of large errors in the vicinity of the singularity, but convergence is still slow.
Figure 7: Example 3.3: for $\beta=4/3$, the approximate solutions $u_{i}$ and basis functions $\varphi_{i}^{NN}$ learned using the standard Galerkin neural network approach in Section 2.1.
@@ 3.1.2 Knowledge-based functions
Suppose $\Omega$ contains a single corner of interest $x^{(0)}$ with interior angle $\alpha^{(0)}$. For instance, the $L$-shaped domain $(-1,1)^{2}\backslash(-1,0)^{2}$ contains a non-convex corner with angle $\alpha^{(0)}=3\pi/2$ at the origin. Solutions of (49) in such a domain have the well-known structure
@@
(53) $\lambda_{j}=\frac{j\pi}{\alpha^{(0)}},\;\;\;j=1,2,3,\dots.$
Figure 8: Example 3.3: for $\beta=4/3$, the approximate solutions $u_{i}$ (top row) and basis functions $\varphi_{i}^{NN}$ learned using the extended Galerkin neural network approach with $\Psi(x;\mu)=r(x)^{\mu}\sin(\mu\theta(x))$ (bottom row). A comparison of the loss function vs. number of learned basis functions is also provided for the basic feedforward network architecture vs. the extended architecture described in Figure 3.
@@
While there is no closed form solution to this problem, we provide the value of the a posteriori error indicator $\langle r(u_{i-1}),\varphi_{i}^{NN}\rangle$ for each basis function learned as well as the basis functions $\varphi_{i}^{NN}$ themselves, which provide a function representation of the error $u-u_{i-1}$. Figure 7 shows the first several basis functions using a standard feedforward neural network for learning each basis function. All hyperparameters for this example are provided in Appendix B. We observe that the approximation converges quite slowly near $x=0$ and the basis functions $\varphi_{i}^{NN}$ show that the magnitude of the pointwise error is large in the vicinity of the singularity.
Next, we repeat the simulation, this time with an augmented neural network architecture. Namely, we compute $\varphi_{i}^{NN}$ according to (29) with $\Psi(r,\theta;\mu)=r^{\mu}\sin(\mu\theta)$. Since the eigenvalues are known, we simply take $\mu_{j}=2j/3$ and $\mathcal{M}=20$. However, we shall demonstrate in Section 4 how Galerkin neural networks may be used to approximate the eigenvalues $\lambda_{j}$ when their values are unknown. Figure 8 shows the analogous results to Figure 7. We observe significantly faster convergence and the approximate pointwise errors decrease steadily from $\mathcal{O}(10^{-2})$ to $\mathcal{O}(10^{-4})$. Additionally, we note that the singularity is clearly captured correctly by the knowledge-based functions.
@@ 3.2.1 Regularity and variational formulation
Most notable is that (54) suggests that the residual of the incompressibility equation in (47) should actually be taken in the $H^{1}_{\beta}$ norm, which leads to the continuous and coercive formulation: seek $(\mathbf{u},p)\in\mathbf{H^{2}_{\beta}}(\Omega)\times H^{1}_{\beta}(\Omega)/\mathbb{R}$ such that
(55) $(-\Delta\mathbf{u}+\nabla p,-&\Delta\mathbf{v}+\nabla q)_{\beta,\Omega}+(\text{div}\;\mathbf{u},\text{div}\;\mathbf{v})_{H^{1}_{\beta}(\Omega)}+\delta(\mathbf{u},\mathbf{v})_{H^{3/2}_{\beta}(\partial\Omega)}\\ &=(\mathbf{f},-\Delta\mathbf{v}+\nabla q)_{\beta,\Omega}+(g,\text{div}\;\mathbf{v})_{H^{1}_{\beta}(\Omega)}+\delta(\mathbf{u}_{D},\mathbf{v})_{\mathbf{H^{3/2}_{\beta}}(\Omega)}\end{aligned}$
for all $(\mathbf{v},q)\in\mathbf{H^{2}_{\beta}}(\Omega)\times H^{1}_{\beta}(\Omega)/\mathbb{R}$. This problem takes the form (13) with $\mathcal{X}:=\mathbf{H^{2}_{\beta}}(\Omega)\times H^{1}_{\beta}(\Omega)/\mathbb{R}$,
(56) $a_{LS}((\mathbf{u},p),(\mathbf{v},q)):=(-\Delta\mathbf{u}$ $+\nabla p,-\Delta\mathbf{v}+\nabla q)_{\beta,\Omega}$
$+(\text{div}\;\mathbf{u},\text{div}\;\mathbf{v})_{H^{1}_{\beta}(\Omega)}+\delta(\mathbf{u},\mathbf{v})_{H^{3/2}_{\beta}(\partial\Omega)},$
and
(57) $F_{LS}((\mathbf{v},q)):=(\mathbf{f},-\Delta\mathbf{v}+\nabla q)_{\beta,\Omega}+(g,\text{div}\;\mathbf{v})_{H^{1}_{\beta}(\Omega)}+\delta(\mathbf{u}_{D},\mathbf{v})_{\mathbf{H^{3/2}_{\beta}}(\Omega)}.$
@@ 3.2.2 Knowledge-based functions
with the constants $A,B,C,D$ chosen so that $\Psi$ satisfies suitable boundary conditions.
Figure 9: Channel with triangular cavity for Example 3.5. The rectangle channel is described by $(-2,2)\times(0,2)$ while the triangular cavity has vertices at $(-1,0)$, $(1,0)$, and $(0,-3)$.
@@
We specify $\mathbf{f}=\mathbf{0}$ and $g=0$. At the reentrant corners, we expect a singularity to manifest in both the velocity and pressure. At the bottom corner, we expect infinitely cascading eddies driven by the channel flow. As described in Section 3.2.2, we shall supplement the standard feedforward network with $\Psi(r,\theta;\mu)=\chi(r)\mathfrak{Re}[r^{\mu}\Psi_{\mu}(\theta)]$ in the bottom corner and $\Psi(r,\theta;\mu)=\chi(r)r^{\mu}\Psi_{\mu}(\theta)$ at the reentrant corners. The exact values of the eigenvalue $\lambda$ are $\lambda=1.58223$ in the reentrant corners and $\lambda=7.56813+3.37941i$ in the bottom corner with $\mathcal{M}=1$. Precise details may be found in Appendix A. As with Example 3.3, we shall demonstrate in Section 4 how these values of $\lambda$ may be approximated using Galerkin neural networks in the event that they are unknown.
Figure 10: Example 3.5: velocity and pressure with $\beta=0$ (top) and $\beta=5/3$ (bottom) in (56) with basis functions learned using the standard Galerkin neural network approach described in Section 2.1. The approximation is non-physical in the channel when the least squares variational problem is posed on unweighted Sobolev spaces. Without enriching the network with knowledge-based functions, the flow behavior at the non-convex corners and in the cavity remains unresolved.
@@
We first demonstrate the importance of choosing the Sobolev weight parameter $\beta$ appropriately. In contrast with Example 3.2, this example contains smooth data $f$ and $g$ and continuous data $\mathbf{u}_{D}$. All hyperparameters for this example are provided in Appendix B. Figure 10 shows the approximated velocities and pressures for $\beta=0$. We shall use $\varphi_{1,i}^{NN}$, $\varphi_{2,i}^{NN}$, and $q_{i}^{NN}$ to denote the $i$th basis functions for the velocities and pressure, respectively. The resulting velocity and pressure fields in Figure 10 show that even after learning six basis functions, the flow is nonphysical; we expect a near constant velocity $u_{1}$ along the $x$-direction in the channel. Figure 10 also shows the analogous results for $\beta=5/3$. Despite not adequately capturing the singularities and the expected sequence of eddies, the approximate solutions do not exhibit spurious oscillations and the behavior in the channel is as expected.
@@
Figure 11: Example 3.5: velocity and pressure with $\beta=5/3$ in (56) and basis functions learned using the extended Galerkin neural network approach with extended neural network architecture (top row) and velocity streamlines with two progressive zooms showing resolution of the Moffatt eddies (bottom row).
@@
We fix $\beta=5/3$ next and approximate the solution using six basis functions learned by extended Galerkin neural networks. We remarkably observe that with the supplemental activation functions, i.e. $\varphi_{1,i}^{NN},\varphi_{2,i}^{NN},q_{i}^{NN}\in V_{\mathbf{n},L}^{\sigma}\oplus V_{\mathcal{M}}^{\Psi}$, both the singular features at the reentrant corners and the eddies in the cavity are captured accurately. As Figure 11 shows, we are able to capture at least ten eddies. The basis functions used to approximate the velocities and pressures are shown in Figure 12.
@@
Finally, when $\beta$ is too large, the solution space contains functions which may have worse singular behavior than the true solution due to the strong damping effect of $r^{\beta}$ on the singularities. In particular, the bilinear operator $a_{LS}$ may admit as arguments functions with far stronger singularities than those of the true solution. We show the results for $\beta=4$ in Figure 13.